Co 38. Does Christine now reject or fail to reject (FTR) Ho and for what reason? FTR because isl<\td b) reject because p-value< α reject because s) reject because p-value >a c) FTR because p-value<a weighs 23 data, and determine the 39. What test do they use and why? a) t-test because its a sample s.d. c) z-test because its a large enough sample d) z-test because its a sample s.d. b) t-test because the sample is small e) none of these 40. The test is not significant at α-.01. However, using the p-value, how confident are we of the test result? a) 95.24% b) 96.88% c) 98.12% d) 99.06% e) none of these When Europeans first came to Australia, they thought that Koalas looked like bears and in fact that they were bears so they were often called Koala bears. Koalas are not bears-they are marsupials and their correct name is Koalas Though often called the koala bear, this cuddly animal is not a bear at all; it is a marsupial, or pouched mammal. After giving birth, a female koala carries her baby in her pouch for about six months wwwtoda
Answer #1

39)

sample size = 23 * 23 = 529

sample size = 529 > 30

by central limit theorem , we can use z-test

option C) is correct

40)

TS= (Xbar - mu)/(Sd/sqrt(n))

= (504.09 - mu)/(40/sqrt(529))

plug the mu from previous question

p-value = 2 P(Z > |TS|)

confidence level = 100 * (1- p-value)